A rectangular vegetable patch has an area of 28 square meters and a perimeter of 22 meters. What are the dimensions of the vegetable patch?
Answer by ewatrrr(24785)

 
Hi,
.
L + w = 11 0r L = 11-w
A = Lw
(11-w)w = 28 m^2
w^2 - 11w + 28 = 0
(w-7)(w-4) = 0
dimensions are: 4m by 7m